OpenStudy (anonymous):

A ball is thrown 20 m/s from the roof of a building 25m high at an angle of 30 degrees below the horizontal. At what height above the ground will the ball strike the side of another building 20m away from the first?

OpenStudy (perl):



OpenStudy (anonymous):

How do you solve for it?

OpenStudy (perl):

the actual path will look like this

OpenStudy (anonymous):

The time a ball takes to hit the building is determined just by its horizontal velocity so you first calculate that. You should get Cos(30)*20 which is 17.3. Then we calculate the time taken to hit the building. Since Speed=Distance/time, Time=Distance/Speed hence Time=20/17.3=1.16 s Now the height is the initial height of the ball minus downward distance traveled. We use the kinematics equation s=ut+0.5*at^2 For initial velocity, do Sin(30)*20=10 Hence s (displacement)=10*1.16+0.5*10*1.16.^2=18.33 m Hence new height=25-18.33=6.67 m Cheers.